OpenStudy (anonymous):

A ship is anchored off a long straight shoreline that runs North and South. From two observation points 18 miles apart on the shore, the bearing of the ships are N31E and S53E. What is the distance from the ship to each of the observation points?

OpenStudy (agl202):

The distances are 8.4394... mi from the northern end and 13.3397 mi from the southern end. This problem is solved by using the sine formula. Call the points A (south), B (north), and S (ship). The angle at S, opposite the 18 mi length is 180 - (53 + 31) = 96 deg. Then 18 / sin 96 = BS / sin 31 = AS / sin 53. Therefore : BS = 8.4394... mi and AS = 13.3397 mi.

OpenStudy (anonymous):

Can you draw a picture of it to help me?

zepdrix (zepdrix):

First off all, the problem should be worded like this: A ship is anchored off a long straight shoreline that runs North and South. From two observation points 18 miles apart on the shore, the bearing of the `ship is` N31E and S53E. What is the distance from the ship to each of the observation points? Lemme draw it a sec, I'm having trouble with this one also.

zepdrix (zepdrix):